How many grams of each of the following substances
Answer #1

2a (1) Ce(IO4)4 :

Solubilty of cerium periodate is 2.72 - 2.88 mol/dm3 = 2.72 - 2.88 mol/litre = 2.72 - 2.88 mol/1000 mL

So for 1000 mL 2.72-2.88 moles can be dissolved

for 100 mL 0.272-0.288 moles i.e 0.272-0.288 moles/100 mL

So 0.288 x M wt /100 mL

0.288 x 839.72 g/100 mL

= 241.83 g/100 mL

so for 100 mL of cold water 241.83 grams can be dissolved.

2. RaSO4

The density of RaSO4 is 5.42 g/cm3.= 5.42 g/mL

So for 100 mL of cold water = 5.42 x 100 = 542 g/100 mL

3. Pb(NO3)2 density = 4.53 g/mL

So for 100 mL of cold water = 4.53 x 100 = 453 g/100 mL

4. (NH4)2SeO4

density = 1.769 g/cm3 (20 °C) = 1.769 g/mL

So for 100 mL of ice cold water = 176.9 g/100 mL can be dissolved.

2b (NH4)2SeO4 is least soluble on gram per 100 mL basis as it has least solubility/100 mL i.e 176.9 g/100 mL

3 Cu2+, Ag+, Ba2+, and Mn2+ & with addition of HBr only Ag+ and Ba+ will precipitate out as follows:

Ag+ + HBr ===> AgBr(s) + H+ On addition of HBr

Ba2+ + H2SO4 ===> BaSO4(s) + 2H+ On addition of H2SO4

Ag+ and Ba+ are present in each of the precipitates. Cu2+ and Mn2+ are not present in the original solution because if present, they would have formed Mn(OH)2(s) and/or Cu(OH)2(s).

4 a molecular: Fe2S3(s) + 6 HBr(aq) → 3 H2S(g) + 2 FeBr3(aq)

total ionic: Fe2S3(s) + 6 H(+) + 6 Br(-) → 3 H2S(g) + 2 Fe(3+) + 6 Br(-)

net ionic: Fe2S3(s) + 6 H(+) → 3 H2S(g) + 2 Fe(3+)

b CuCO3 is an insoluble salt so the ionic equation is

Molecular equation:  Cu(NO3)2 (aq) + K2CO3 (aq) --> CuCO3 (s) + 2KNO3 (aq)

Total ionic equation: Cu2+ (aq)+ 2 NO3- (aq)+ 2 K+ (aq)+ CO32-(aq) = CuCO3 (s)+ 2 K+ (aq) + 2 NO3- (aq)
  
Net ionic equation: Cu2+ (aq) + CO32- (aq)= CuCO3 (s)

c No reaction because there is a reaction only if a product is insoluble, a gas, or a nonelectrolyte but all the above substances are soluble electrolytes, so there is no reaction.

d Molecular equation: HNO3 (aq) + Bi(OH)3 (s) ----> Bi(NO3)3 (aq) + 3 H2O (l)

Net Ionic Equation : 3 [ H+ ] (aq) + Bi(OH)3 (s) ----> [ Bi3+ ] (aq) + 3 H2O (l)